# Rethinking Progression of Memory State in Robotic Manipulation: An Object-Centric Perspective

### v1

@@ Abstract
Techical, Data & Code Appendix:https://libero-mem.github.io
@@ Introduction
Figure 1: LIBERO-Mem: robotic manipulation tasks of object-level PO-NM dependencies. These tasks require memory of prior actions and object-specific state tracking beyond what purely Markovian or fully observable policies can handle, highlighting the importance of persistent, object-specific memory for short- and long-horizon reasoning across visually similar inputs. In (a) OM, the robot must recall its last action (e.g., pick up or place down) to act correctly. In (b) OS, success depends on remembering how many times an object has been manipulated, since visual cues are insufficient. In (c) OR, the robot must track the temporal order of object relations and interactions (e.g., from left to right). In (d) $O$, occluded objects require the robot to rely on memory of past placements to identify targets.
@@
Humans effortlessly recall past interactions with specific objects, such as where they last placed a salt bottle or whether they have already sprinkled salt into a pot of soup a number of times, enabling them to carry out long-horizon, multiple-step tasks with precision. This object-level memory plays a vital role for avoiding redundant actions or missing steps, especially in tasks involving repetitive steps, visually similar items, and extended temporal dependencies. Such object-level challenges are inherently partially observable and non-Markovian (PO-NM) (illustrated in Fig. 1), where the agent’s current observation does not fully capture the true state of the environment, and optimal decisions depend on the history of interactions associated with specific objects.
@@
Conversely, robotic visuomotor policies typically rely solely on the most recent sensory inputs to determine subsequent actions  (DBLP:OpenVLA2024; DBLP:RoboFlamingo2024; DBLP:RoboAgent_2024; DBLP:RT12023; DBLP:Octo2024; zawalski2024robotic; wang2024hpt; DBLP:conf/rss/YangGBSVFSL24; tian2024tokenize), lacking mechanisms to encode and recall object-centric history (BridgeDatav2_2023; DBLP:RTX2024; DROID2024; james2020rlbench; libero2023; robocasa2024). Likewise, most robotic benchmarks (james2020rlbench; libero2023; robocasa2024; mu2021maniskill; gu2023maniskill2; DBLP:journals/corr/maniskill3_2024) primarily focus on evaluating policies in short-horizon or atomic tasks. Although several benchmarks have emerged to evaluate long-horizon tasks that requires memory modeling (memoryBench2025; MIKASArobo2025; morad2023popgym; chevalier2023minigrid; pleines2025memory), they largely overlook the challenges posed by PO-NM settings.
@@
To address this gap in benchmarking, we introduce LIBERO-Mem, a new suite of manipulation tasks specifically designed to evaluate a model’s ability to retain object-centric interactions over time. Unlike prior memory benchmarks, LIBERO-Mem emphasizes object-level memory under ambiguity in object identity, location, and relational history, explicitly targeting complex PO-NM scenarios. The suite includes four task types: (a) object motion (OM), (b) object sequence (OS), (c) multi-object sequence (OR), and (d) multi-object occlusion (OO). Each task type is designed to evaluate different aspects of transient and persistent memory by introducing ambiguities that can only be resolved through temporal reasoning over the history of object interactions. By focusing on memory under uncertainty, LIBERO-Mem takes a key step toward enabling general-purpose robots to operate reliably in everyday settings, where even seemingly simple tasks, such as identifying the correct object to manipulate, become challenging without awareness of prior interactions.
@@
As a step towards overcoming the limitations of existing visuomotor policies, we propose Embodied-SlotSSM, a novel VLA framework that integrates slot-based state-space modeling (SSM) (jiang2024slot) to maintain structured, object-centric memory over time. Embodied-SlotSSM encodes temporal dynamics and inter-object interactions into discrete, persistent slots, enabling consistent tracking of entities throughout a task. This structured memory supports robust state estimation and informed decision-making, particularly in partially observable and non-Markovian settings where access to interaction history is critical. By grounding both perception and action in visual semantics and high-level task intent, E-SlotSSM enables more adaptive and reliable behavior under real-world uncertainty.
@@
To highlight the limitations of existing works and the challenges posed by our benchmark, we evaluate Embodied-SlotSSM and prior state-of-the-art VLA models on the benchmark of the LIBERO-Goal (libero2023), and our proposed LIBERO-Mem benchmark for PO-NM tasks. We also carry out studies of our proposed method in capturing long-term dependencies, improving action prediction, and enhancing manipulation efficiency. Our results highlight the practicality and scalability of slot-based memory representations, paving the way for more reliable and memory-efficient robotic systems in non-Markovian environments.
@@
  • We conduct experiments in both general Markovian and special non-Markovian settings, showing that Embodied-SlotSSM enhances stateful reasoning, long-horizon action prediction, and performance on manipulation tasks.
@@
Task designs for object-centric PO-NM. LIBERO-Mem consists of 10 tasks spanning four object-centric memory dimensions: Object Motion (OM), Object Sequence (OS), Object Relations (OR), and Object Occlusion (OO). Each task presents temporal dependencies and ambiguity requiring structured memory beyond instantaneous observations. Formally, let $\mathcal{T}_{i}=\{o_{t}^{i}\}_{t=1}^{T}$ denote a trajectory for task $i$, where $o_{t}^{i}$ is the visual observation at time $t$. For each task $i$, there exist at least two trajectories $\mathcal{T}_{i}^{(1)}$, $\mathcal{T}_{i}^{(2)}$ such that $\exists\,t_{1},t_{2}$ with $o_{t_{1}}^{(1)}=o_{t_{2}}^{(2)}$, yet their underlying task states or required actions differ. This guarantees that visual observations are not uniquely predictive of the correct behavior, thereby enforcing the need for structured memory across time.
@@
As the underlying process of visual representations is inherently modular, SlotSSM (jiang2024slot) is coupled with a slot encoder (locatello2020object) to decompose $\mathbf{e}_{t}$ into $K$ individual slot representations $\mathbf{e}_{t}=\left[\mathbf{s}_{t},...,\mathbf{s}_{t}\right]$, thereby maintaining $\mathbf{h}_{t}$, $\mathbf{y}_{t}$ also modularly as $\mathbf{h}_{t}=\text{concat}\left[\mathbf{h}^{1}_{t},...,\mathbf{h}^{K}_{t}\right]$, and $\mathbf{y}_{t}=\text{concat}\left[\mathbf{y}^{1}_{t},...,\mathbf{y}^{K}_{t}\right]$. Thus, the matrices $\overline{A}_{t}$, $\overline{B}_{t}$, and $C_{t}$ are designed to be block-diagonal, where each block is conditioned solely on the corresponding slot input.
@@
Slot Attention for Object Localization
Manipulation requires isolating and tracking discrete object entities, yet conventional visual encoders often produce entangled representations that conflate background and object-level signals. To address this, we apply Slot Attention (locatello2020object) as a tokenization function $g(\cdot)$ that transforms dense visual embeddings $\mathbf{v}_{t}\in\mathbb{R}^{K\times D_{\mathtt{enc}}}$ into a set of modular object-centric tokens $\mathbf{s}_{t}=\{\mathbf{s}_{t}^{1},\ldots,\mathbf{s}_{t}^{N}\}$, where $\mathbf{s}_{t}\in\mathbb{R}^{N\times D_{\mathtt{slot}}}$. Slot Attention iteratively binds spatial feature patches to a fixed number of learnable object queries via attention and recurrent updates, yielding disentangled representations that reflect the semantics of $\mathbf{v}_{t}$ (xu2022groupvit; DBLP:conf/iclr/JiaLH23). The slot update at time $t$ is computed via multi-head attention followed by a recurrent refinement:
@@
Then, to condition action generation on both the structured memory and the current scene, we introduce a lightweight RelationEncoder that produces 16 relational tokens. This module performs cross-attention between the slot latents ${\{\mathbf{d}_{t}^{(j)}}\}_{j=1}^{K}$ and raw visual features $\mathbf{v}_{t}$ to produce $L$ relation tokens ${\{\mathbf{r}_{t}^{(j)}}\}_{j=1}^{L}$, enabling context-aware reasoning over object states and interactions. The final action $\hat{\mathbf{a}}_{t}$ is decoded through a VLA head conditioned on: (1) the relation tokens $\{\mathbf{r}_{t}^{(j)}\}_{j=1}^{L}$, (2) the slot dynamics $\{\mathbf{d}_{t}^{(j)}\}_{j=1}^{K}$, and (3) the current task query embedding $l$:
@@
We evaluate our model on object-centric manipulation tasks using the LIBERO-Goal and the LIBERO-Mem benchmark to respectively evaluate general and object-level PO-NM task performance. All experiments are conducted in simulation with fixed scene layouts and randomly initialized object poses. Each task is repeated for $N$ multiple seeds, where action accuracy is computed as the ratio of successful completions over total attempts ($\text{success rate}=\frac{\text{n${}_{0}$ success}}{N}$) in Table 3, or subgoal completion $\frac{\text{subgoals completed}}{\text{total subgoals}}$ over $N$ seeds in Table 4. We evaluate comparatively against Naive E-SlotSSM using OpenVLA (DBLP:OpenVLA2024) with slot attention (locatello2020object) (as object-centric version of SlotVLA (slotvla2025)), $\pi_{0}$ (black2024pi0visionlanguageactionflowmodel), each with horizon $h$ denoting number of input frames, over $N=20$.
@@
Our empirical Naive E-SlotSSM achieves the highest success rates on LIBERO-Goal, outperforming slot-based baselines. The results in Table 3 highlight the effectiveness of structured object-centric memory in non-Markovian manipulation tasks. While SlotVLA ($h=8$) achieves moderate performance ($66.5\%$) by extending temporal context length, it still struggles on tasks involving multi-object interactions and occlusions (e.g., middle drawer open, top drawer open $\rightarrow$ bowl in). In contrast, Naive E-SlotSSM consistently outperforms others ($\mathbf{80.1\%}$ avg), demonstrating robustness across both simple and temporally entangled tasks. Notably, it handles action sequences involving relational reasoning (e.g., bottle in rack) and spatial displacements (e.g., bowl in plate) reliably, attributed to its ability to integrate persistent slot tracking memory. These results affirm the usefulness of memory design for generalizing over object-centric tasks in general manipulation.
Figure 3: Slot-attention visualization in task T1: gripper and bowl slots (bbox, attention) as robot lifts and places the bowl.
Object PO-NM Task Performance with LIBERO-Mem:
@@ Conclusion
Limitations: LIBERO-Mem is a simulated setting for future physical extension. The empirical version of Embodied-SlotSSM (Naive E-SlotSSM) currently serves as a weak baseline that leverages oracle subgoal representations rather than inferring it autonomously, leaving open the challenge of self-discovered subgoal reasoning and PO-NM at object level.

### v2

@@ Abstract
Extended Versionhttps://libero-mem.github.io
@@ Introduction
Figure 1: LIBERO-Mem: robotic manipulation tasks of object-level POMDP dependencies. These tasks require memory of prior actions and object-specific state tracking beyond what purely Markovian or fully observable policies can handle, highlighting the importance of persistent, object-specific memory for short- and long-horizon reasoning across visually similar inputs. In (a) object motion (OM), the robot must recall its last action (e.g., pick up or place down) to act correctly. In (b) object sequence (OS), success depends on remembering how many times an object has been manipulated, since visual cues are insufficient. In (c) multi-object sequence (OR), the robot must track the temporal order of object relations and interactions (e.g., from left to right). In (d) multi-object occlusion ($O$), occluded objects require the robot to rely on memory of past placements to identify targets.
@@
Humans effortlessly recall past interactions with specific objects, such as where they last placed a salt bottle or whether they have already sprinkled salt into a pot of soup a number of times, enabling them to carry out long-horizon, multiple-step tasks with precision. This object-level memory plays a vital role in avoiding redundant actions or missing steps, especially in tasks involving repetitive steps, visually similar items, and extended temporal dependencies. Such object-level challenges are inherently partially observable Markov decision processes (POMDP) (illustrated in Fig. 1), where the agent’s current observation does not fully capture the true state of the environment, and optimal decisions depend on the history of interactions associated with specific objects.
@@
Conversely, robotic visuomotor policies typically rely solely on the most recent sensory inputs to determine subsequent actions  (DBLP:OpenVLA2024; DBLP:RoboFlamingo2024; DBLP:RoboAgent_2024; DBLP:RT12023; DBLP:Octo2024; zawalski2024robotic; wang2024hpt; DBLP:conf/rss/YangGBSVFSL24; tian2024tokenize), lacking mechanisms to encode and recall object-centric history (BridgeDatav2_2023; DBLP:RTX2024; DROID2024; james2020rlbench; libero2023; robocasa2024). Likewise, most robotic benchmarks (james2020rlbench; libero2023; robocasa2024; mu2021maniskill; gu2023maniskill2; DBLP:journals/corr/maniskill3_2024) primarily focus on evaluating policies in short-horizon or atomic tasks. Although several benchmarks have emerged to evaluate long-horizon tasks that require memory modeling (memoryBench2025; MIKASArobo2025; morad2023popgym; chevalier2023minigrid; pleines2025memory), they largely overlook the challenges posed by POMDP settings.
@@
To address this gap in benchmarking, we introduce LIBERO-Mem, a new suite of manipulation tasks specifically designed to evaluate a model’s ability to retain object-centric interactions over time. Unlike prior memory benchmarks, LIBERO-Mem emphasizes object-level memory under ambiguity in object identity, location, and relational history, explicitly targeting complex POMDP scenarios. The suite includes four task types: (a) object motion (OM), (b) object sequence (OS), (c) multi-object sequence (OR), and (d) multi-object occlusion (OO). Each task type is designed to evaluate different aspects of transient and persistent memory by introducing ambiguities that can only be resolved through temporal reasoning over the history of object interactions. By focusing on memory under uncertainty, LIBERO-Mem takes a key step toward enabling general-purpose robots to operate reliably in everyday settings, where even seemingly simple tasks, such as identifying the correct object to manipulate, become challenging without awareness of prior interactions.
@@
As a step towards overcoming the limitations of existing visuomotor policies, we propose Embodied-SlotSSM, a novel VLA framework that integrates slot-based state-space modeling (SSM) (jiang2024slot) to maintain structured, object-centric memory over time. Embodied-SlotSSM encodes temporal dynamics and inter-object interactions into discrete, persistent slots, enabling consistent tracking of entities throughout a task. This structured memory supports robust state estimation and informed decision-making, particularly in partially observable and non-Markovian settings where access to interaction history is critical. By grounding both perception and action in visual semantics and high-level task intent, Embodied-SlotSSM enables more adaptive and reliable behavior under real-world uncertainty.
@@
To highlight the limitations of existing works and the challenges posed by our benchmark, we evaluate Embodied-SlotSSM and prior state-of-the-art VLA models on the benchmark of the LIBERO-Goal (libero2023), and our proposed LIBERO-Mem benchmark for POMDP tasks. We also carry out studies of our proposed method in capturing long-term dependencies, improving action prediction, and enhancing manipulation efficiency. Our results highlight the practicality and scalability of slot-based memory representations, paving the way for more reliable and memory-efficient robotic systems in non-Markovian environments.
@@
  • We conduct experiments in both general Markovian and special non-Markovian settings via an oracle-supported implementation of Embodied-SlotSSM, denoted Naive E-SlotSSM, showing that Embodied-SlotSSM enhances stateful reasoning, long-horizon action prediction, and performance on manipulation tasks.
@@
Task designs for object-centric POMDP. LIBERO-Mem consists of 10 tasks spanning four object-centric memory dimensions: Object Motion (OM), Object Sequence (OS), Object Relations (OR), and Object Occlusion (OO). Each task presents temporal dependencies and ambiguity requiring structured memory beyond instantaneous observations. Formally, let $\mathcal{T}_{i}=\{o_{t}^{i}\}_{t=1}^{T}$ denote a trajectory for task $i$, where $o_{t}^{i}$ is the visual observation at time $t$. For each task $i$, there exist at least two trajectories $\mathcal{T}_{i}^{(1)}$, $\mathcal{T}_{i}^{(2)}$ such that $\exists\,t_{1},t_{2}$ with $o_{t_{1}}^{(1)}=o_{t_{2}}^{(2)}$, yet their underlying task states or required actions differ. This guarantees that visual observations are not uniquely predictive of the correct behavior, thereby enforcing the need for structured memory across time.
@@
As the underlying process of visual representations is inherently modular, SlotSSM (jiang2024slot) is coupled with a slot encoder (locatello2020object) to decompose $\mathbf{e}_{t}$ into $K$ individual slot representations $\mathbf{e}_{t}=\left[\mathbf{s}^{1}_{t},...,\mathbf{s}^{K}_{t}\right]$, thereby maintaining $\mathbf{h}_{t}$, $\mathbf{y}_{t}$ also modularly as $\mathbf{h}_{t}=\text{concat}\left[\mathbf{h}^{1}_{t},...,\mathbf{h}^{K}_{t}\right]$, and $\mathbf{y}_{t}=\text{concat}\left[\mathbf{y}^{1}_{t},...,\mathbf{y}^{K}_{t}\right]$. Thus, the matrices $\overline{A}_{t}$, $\overline{B}_{t}$, and $C_{t}$ are designed to be block-diagonal, where each block is conditioned solely on the corresponding slot input. In our work, $K=16$ by default.
@@
Slot Attention for Object Localization.
Manipulation requires disentangling and tracking discrete object entities, yet conventional visual encoders often produce entangled representations that conflate background and object-level signals. To address this, we apply Slot Attention (locatello2020object) as a tokenization function $g(\cdot)$ that transforms dense visual embeddings $\mathbf{v}_{t}\in\mathbb{R}^{K\times D_{\mathtt{enc}}}$ into a set of modular object-centric tokens $\mathbf{s}_{t}=\{\mathbf{s}_{t}^{1},\ldots,\mathbf{s}_{t}^{N}\}$, where $\mathbf{s}_{t}\in\mathbb{R}^{N\times D_{\mathtt{slot}}}$. Slot Attention iteratively binds spatial feature patches to a fixed number of learnable object queries via attention and recurrent updates, yielding disentangled representations that reflect the semantics of $\mathbf{v}_{t}$ (xu2022groupvit; DBLP:conf/iclr/JiaLH23). The slot update at time $t$ is computed via multi-head attention followed by a recurrent refinement:
@@
Then, to condition action generation on both the structured memory and the current scene, we introduce a lightweight Relation Encoder that produces 16 relational tokens (for $K=16$ slots). This module performs cross-attention between the slot latents ${\{\mathbf{d}_{t}^{(j)}}\}_{j=1}^{K}$ and raw visual features $\mathbf{v}_{t}$ to produce $L$ relation tokens ${\{\mathbf{r}_{t}^{(j)}}\}_{j=1}^{L}$, enabling context-aware reasoning over object states and interactions. The final action $\hat{\mathbf{a}}_{t}$ is decoded through a VLA head conditioned on: (1) the relation tokens $\{\mathbf{r}_{t}^{(j)}\}_{j=1}^{L}$, (2) the slot dynamics $\{\mathbf{d}_{t}^{(j)}\}_{j=1}^{K}$, and (3) the current task query embedding $l$:
@@
We evaluate our model on object-centric manipulation tasks using the LIBERO-Goal and the LIBERO-Mem benchmark to respectively evaluate general and object-level POMDP task performance. All experiments are conducted in simulation with fixed scene layouts and randomly initialized object poses. Each task is repeated for $N$ multiple seeds, where action accuracy is computed as the ratio of successful completions over total attempts ($\text{success rate}=\frac{\text{n${}_{0}$ success}}{N}$) in Table 3, or subgoal completion $\frac{\text{subgoals completed}}{\text{total subgoals}}$ over $N$ seeds in Table 4. We evaluate comparatively against Naive E-SlotSSM using OpenVLA (DBLP:OpenVLA2024) with slot attention (locatello2020object) (as object-centric version of SlotVLA (slotvla2025)), $\pi_{0}$ (black2024pi0visionlanguageactionflowmodel), each with horizon $h$ denoting number of input frames, over $N=20$.
@@
Our empirical Naive E-SlotSSM achieves the highest success rates on LIBERO-Goal, outperforming slot-based baselines. The results in Table 3 highlight the effectiveness of structured object-centric memory in non-Markovian manipulation tasks. While SlotVLA ($h=8$) achieves moderate performance ($75.5\%$) by extending temporal context length, it still struggles on tasks involving multi-object interactions and occlusions (e.g., middle drawer open, top drawer open $\rightarrow$ bowl in). In contrast, Naive E-SlotSSM consistently outperforms others ($\mathbf{83.0\%}$ avg), demonstrating robustness across both simple and temporally entangled tasks. Notably, it handles action sequences involving relational reasoning (e.g., bottle in rack) and spatial displacements (e.g., bowl in plate) reliably, attributed to its ability to integrate persistent slot tracking memory. These results affirm the usefulness of memory design for generalizing over object-centric tasks in general manipulation.
Figure 3: Slot visualization in task T1: gripper and bowl slots (bbox, attention) as robot lifts and places the bowl down.
Object POMDP Task Performance with LIBERO-Mem:
@@ Conclusion
Limitations: LIBERO-Mem is a simulated setting for future physical extension. The empirical version of Embodied-SlotSSM (Naive E-SlotSSM) currently serves as a weak baseline that leverages oracle subgoal representations rather than inferring it autonomously, leaving open the challenge of self-discovered subgoal reasoning and POMDP at object level.
@@
Appendix
We provide additional details on dataset usage and computational experiments to support reproducibility of our LIBERO-Mem benchmark and Embodied-SlotSSM framework. Readers are encouraged to check our extended version for any updates, clarifications, and additional analyses released after publication.
Dataset Usage
Motivation for Dataset Selection.
We evaluate our method on two benchmark families: the original LIBERO-Goal benchmark and our proposed LIBERO-Mem suite. LIBERO-Goal (libero2023) contains Markovian manipulation tasks and serves as a baseline for standard visuomotor performance. LIBERO-Mem introduces non-Markovian, memory-centric tasks that stress-test object permanence, temporal dependencies, relational reasoning, and occlusion. These datasets together enable evaluating how Embodied-SlotSSM handles both conventional and memory-dependent visuomotor challenges.
Novel Dataset Components.
LIBERO-Mem will be released publicly at a public link  [^1] https://libero-mem.github.io and includes ten newly constructed tasks across four memory dimensions (Object Motion, Object Sequence, Object Relations, and Object Occlusion). Each task is provided with metadata describing temporal dependencies, subgoal structures, and expected memory behaviors. We provide analyses of the dataset in Fig. 1 and Fig. 2. Furthermore, we show examples of OM, OS, OR, OO settings in Fig. 3, Fig. 4, Fig. 5, Fig. 6 with RGB image longside object boxes, object memory, object depth and object mask.
Figure 1: The histogram of frame count across different subsets of LIBERO and our proposed LIBERO-Mem.
Figure 2: The detailed histograms of frame count across different subsets of LIBERO and our proposed LIBERO-Mem.
Public Release of Novel Datasets.
We release LIBERO-Mem with: RGB frames, object masks, subgoal labels, trajectory metadata, and task specifications. The dataset will be accessible under a research-friendly license.  [^2] https://huggingface.co/datasets/libero-mem/LIBERO-Mem
Existing Datasets and Citations.
All datasets drawn from existing literature, particularly LIBERO-Goal (libero2023), are properly cited. We follow the data setting used in OpenVLA  [^3] https://github.com/moojink/rlds_dataset_builder in data building.
Public Availability.
All datasets used in this study are publicly available or will be made public at the links provided. No proprietary or restricted-access data are used.
Non-Public Data.
There are no non-public datasets involved in the experiments presented in this paper.
Computational Experiments
In this section, we discuss the implementation of Naive E-SlotSSM as a representation for Embodied-SlotSSM,
Data Pre-processing Code.
We release the full pre-processing pipeline, including scripts for extracting RGB observations, aligning instance masks, generating slot tokens, and constructing subgoal labels. All code paths needed for reproducing LIBERO-Mem training inputs are documented in the codebase.
Experimental Source Code.
The Naive Embodied-SlotSSM implementation are to be released publicly [^4] https://github.com/libero-mem/naive-e-slotssm and the evaluation toolkit for LIBERO-Mem are also to be released publicly [^5] https://github.com/libero-mem/libero-mem. Here, we also show that Naive Embodied-SlotSSM can systematically addresses the token scaling problems when memory grounding is important, as shown in Fig. 7. The evaluation algorithm is shown in Algorithm 1, and the difference between our evaluation strategy versus an existing one is shown in Table 1. In order to perform evaluation, our subgoals are exemplified as follows,
OM Task: pick up bowl and place it back on the plate 1 time (:goal (Sequence (And (On akita_black_bowl_1 plate_1)) ))
OS Task: pick up bowl and place it back on the plate 3 times (:goal (Sequence (And (On akita_black_bowl_1 plate_1)) (And (On akita_black_bowl_1 plate_1)) (And (On akita_black_bowl_1 plate_1)) ))
OR Task: swap the 2 bowls on their plates using the empty plate (:goal (Or (Sequence (And (On akita_black_bowl_1 plate_3)) (And (On akita_black_bowl_2 plate_1)) (And (On akita_black_bowl_1 plate_2)) ) (Sequence (And (On akita_black_bowl_2 plate_3)) (And (On akita_black_bowl_1 plate_2)) (And (On akita_black_bowl_2 plate_1)) )))
$O$ Task: put the cream cheese in the nearest basket and place that basket in the center (:goal (Or (Sequence (And (In cream_cheese_1 basket_1_contain_region)) (And (On basket_1 kitchen_table_the_center)) ) (Sequence (And (In cream_cheese_1 basket_2_contain_region)) (And (On basket_2 kitchen_table_the_center)) )))
$O$ Task: put the cream cheese in the nearest basket and place the empty basket in the center (:goal (Or (Sequence (And (In cream_cheese_1 basket_1_contain_region)) (And (On basket_2 kitchen_table_the_center)) ) (Sequence (And (In cream_cheese_1 basket_2_contain_region)) (And (On basket_1 kitchen_table_the_center)) )))
If after the last subgoal the robot again changes the goal state, then it means the task has failed by over-repetition, scoring.
Figure 3: Example of OM setting with T1.
Figure 4: Example of OS setting with T3.
Figure 5: Example of OR setting with T8.
Figure 6: Example of $O$ setting with T9.
Figure 7: Token scaling challenges under different temporal window sizes vs visual token count (left) and naive performances qualitatively shown with our simple OM task of pick and place down (right). Slot refers to use of SlotVLA (slotvla2025), and Dense refers to use of OpenVLA (DBLP:OpenVLA2024). We found that concatenation-based methods struggles to pick up and place down, either stuck when it needs to go up or stuck when it needs to go down, probably during training, the model is confused by same visual states but different directions. Meanwhile, our method could pick up the bowl and plate on the plate.
Table 1: Comparison between existing LIBERO goal specification and our compositional extension.
Aspect Existing LIBERO Ours
Goal Type Flat conjunction (And) Composed logic (Sequence, Or, nested And)
Place bowl on plate
Difficulty Level Low-Med: one-/two-step goals Low-High: multi-step, multi-object goals
Rearranging objects across locations
[…]
Place object on 3 plates in order
[…]
Two valid sequences for table setup
[…]
Ordered subgoal execution
Algorithm 1 High-level hierarchical goal evaluation
1:def check_success(goal_state, X):
2:  # X: dict of object states
3:  if is_structured(goal_state):
4:   return eval_goal(goal_state, X, satisfied=[])
5:  else:
6:   return eval_conjunction(goal_state, X)
7:
8:def eval_goal(G, X, satisfied):
9:  op = G[0]
10:  if op == "and":
11:   return eval_conjunction(G[1:], X)
12:
13:  elif op == "sequence":
14:   k = len(satisfied)
15:   if k >= len(G) - 1:
16:    return True
17:   sub = G[k+1] # next subgoal
18:   if eval_goal(sub, X, satisfied):
19:    satisfied.append(sub)
20:   return (len(satisfied) >= len(G) - 1)
21:
22:  elif op == "or":
23:   for subseq in G[1:]:
24:    if not is_prefix_consistent(satisfied, subseq[1:]):
25:     continue
26:    if eval_goal(subseq, X, satisfied.copy()):
27:     return True
28:   return False
29:
30:  else:
31:   raise ValueError("unknown operator")
32:
33:def eval_conjunction(states, X):
34:  return all(eval_predicate(s, X) for s in states)
35:
36:def eval_predicate(state, X):
37:  if len(state) == 3:
38:   p, o1, o2 = state
39:   return eval_pred_fn(p, X[o1], X[o2])
40:  else:
41:   p, o = state
42:   return eval_pred_fn(p, X[o])
Public Release of Code.
Upon publication, the entire codebase (data loaders, training pipelines, baseline implementations, and evaluation utilities) will remain publicly available under a research-permissive license.
Implementation Comments and Documentation.
Newly introduced components (Slot Dynamics Encoder, windowed SlotSSM predictor, slot fusion module, relation encoder, and VLA decoding head) include comments referencing their mathematical formulations. These comments are designed to support line-by-line tracing of each algorithmic step.
Randomness and Seed Control.
For experiments with stochastic behavior, we set global seeds for Python, NumPy, and PyTorch. Unless otherwise stated, each result averages over $N=20$ runs with distinct seeds.
Computing Infrastructure.
The models are finetuned on 2x NVIDIA A100 GPUs. To support reproducibility, the complete list of packages are included in our requirements.txt file included in the code repository. Readers may refer to the extended version for any additional environment details released after publication.
Evaluation Metrics.
We report: (1) Task Success Rate: full completion of the LIBERO instruction. (2) Subgoal Completion Rate: symbolic subgoal satisfaction using Sequence ($\rightarrow$) and Or ($\lor$) operators for LIBERO-Mem tasks. Further details of how these subgoals are set up are also provided below. These metrics capture both overall task performance and fine-grained memory-sensitive progress.
Number of Runs.
Each experimental task configuration is evaluated over $20$ independent rollouts with different seeds.
Variation and Distributional Measures.
We report means across seeds per task or average subgoal completion in the main text. In the extended version, we aim to additionally provide standard deviations and confidence intervals to characterize performance stability.
Final Hyperparameters.
This accepted version does not include an extended hyperparameter search. Instead, we adopt default hyperparameters from the originally cited papers. Specifically,
  • We use $K=16$ slots to reduce the chance of missing small or background objects.
  • The temporal prediction horizon is fixed to $P=32$ frames, matching the observation chunk size during training.
  • All other hyperparameters follow the defaults from prior work unless explicitly noted.
Hyperparameter Search Protocol.
Sweep-based hyperparameter search was limited as all hyperparameters are fixed from defaults or prior literature. Evaluation is conducted on a representative subset of LIBERO-Mem covering all four memory dimensions.